Determine the in-order, pre-order, and post-order sequence of the following: 2 7 5 2 6 9 5 11 4 30 (20 (39 10 25 35 (42 15 23


1
Expert's answer
2021-06-10T05:16:27-0400

Strange way to determine the tree. I printed in-order, pre-order, and post-order sequence of the "2 7 5 2 6 9 5 11 4 30". You can change it after the comment "Determinate graph". Good luck!

class Node:
    def __init__(self, key):
        self.left = None
        self.right = None
        self.val = key


def printInorder(root):
    if root:
        printInorder(root.left)
        print(root.val),
        printInorder(root.right)


def printPostorder(root):
    if root:
        printPostorder(root.left)
        printPostorder(root.right)
        print(root.val),


def printPreorder(root):
    if root:
        print(root.val),
        printPreorder(root.left)
        printPreorder(root.right)


# Determinate graph ------> 2 7 5 2 6 9 5 11 4 30
root = Node(2)
root.left = Node(7)
root.right = Node(5)
root.left.left = Node(2)
root.left.right = Node(6)
root.right.left = Node(9)
root.right.right = Node(5)
root.left.left.left = Node(11)
root.left.left.right = Node(4)
root.left.right.left = Node(30)


print("Preorder traversal of binary tree is")
printPreorder(root)
print("\nInorder traversal of binary tree is")
printInorder(root)
print("\nPostorder traversal of binary tree is")
printPostorder(root)